Prove the identity: sin^2 theta/1-cos theta = (sec theta + 1)/sec theta
Answer by stanbon(75887)

Prove the identity: sin^2 theta/1-cos theta = (sec theta + 1)/sec theta
------------
sin^2/(1-cos) = (sec+1)/sec
-----
Cross multiply to get:
-----
sin^2*sec = (1-cos)(1+sec)
---
sin^2/cos = 1+sec-cos-cos*sec
---
sin^2/cos = sec-cos +1-1
---
Multiply both sides by cos to get:
sin^2 = 1-cos^2
----
sin^2 = sin^2
=========================
Cheers,
Stan H.